OpenStudy (anonymous):

write the expression as a single logarithm 4 log^by +3log^bx HELP Please

OpenStudy (anonymous):

@amoodarya can you help with this one

OpenStudy (amoodarya):

b is base of log ?

OpenStudy (anonymous):

\[\ln a + \ln b = \ln ab\]\[b \ln a = \ln a^b\]

OpenStudy (amoodarya):

\[4 \log_{b} y +3 \log_{b} x = \log_{b} y^{4} + \log_{b} x^{3} = \log_{b} y^{4} x^{3}\]